The formula for the perimeter of a rectangle is P=2L+2w, where L is the length and w is the width. A rectangle has a perimeter of 24 inches. Find its dimensions if its length is 3 inches greater than its width.
Answer by checkley77(12844)

L=W+3
PERIMETER=2L+2W
24=2(W+3)+2W
24=2W+6+2W
24=4W+6
4W=24-6
4W=18
W=18/4
W=4.5 ANS. FOR THE WIDTH.
L=4.5+3=7.5 ANS. FOR THE LENGTH.
PROOF:
24=2*7.5+2*4.5
24=15+9
24=24